# TALM: Dynamic Tree-Structured Multi-Agent Framework with Long-Term Memory for Scalable Code Generation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
We evaluate TALM on HumanEval (Chen et al., 2021), BigCodeBench (Zhuo et al., 2025), and ClassEval (Du et al., 2023) benchmarks. Results show that TALM outperforms all state-of-the-art approaches while maintaining high token efficiency. Beyond benchmark results, TALM provides a general paradigm for integrating structured reasoning with long-term memory, suggesting a path forward for scalable, reliable and self-improving code generation agents.
@@ 2. Related Work
AutoGen (Wu et al., 2023) provides a general multi-agent platform where a centralized planner coordinates multiple specialized agents through role-playing and structured dialogues. This design improves modularity and allows flexible integration of external tools TDAG (Wang et al., 2025) dynamically decomposes tasks into executable steps and instantiates specialized sub-agents as needed. Its reflection mechanism enables feedback and backtracking, enhancing robustness and recovery. AgentNet (Yang et al., 2025) employs a dynamically evolving DAG communication topology, allowing agents to adapt roles and strategies through an evolution phase that fosters specialization and autonomous task allocation. In contrast, ChatDev (Qian et al., 2024) focuses on simulating the software development life cycle by assigning agents to roles such as analyst, designer, coder, and tester, with dedicated mechanisms for cross-stage context menagement. MetaGPT (Hong et al., 2024) emphasizes responsibility separation: each agent operates within well-defined input–output boundaries, and communication is achieved through a subscription-based messaging mechanism.
@@ 3. TALM Framework
Figure 1. TALM workflow overview. The root agent decomposes the task into a tree of subtasks; child agents solve localized steps and return results upward.:
@@ 3.1. Tree-Structured Collaboration
(a) Child-Agent Clarification (bottom-up).:
(b) Structure Correction (top-down).:
Figure 2. Localized re-reasoning in TALM. Left: child-to-parent clarification when task specs are underspecified; Right: parent-initiated structure correction when decomposition is flawed.:
To prevent uncontrolled tree growth, TALM introduces three hyperparameters: the maximum depth $m$, the initial branching factor $n$, and the decay rate $k$ ($m,n,k\in\mathbb{Z}^{+}$). A node may create children only if its depth is less than $m$, ensuring that nodes at depth $m$ are leaves. The number of children a node may spawn begins at $n$ and decreases by $k$ at each subsequent level, producing a narrowing tree shape. This decay mechanism encourages deeper nodes to concentrate on highly specific subtasks, while also constraining overall complexity. Together, these parameters regulate the height and width of the collaboration tree, striking a balance between flexibility, resource efficiency, and structural controllability.
@@ 3.2. Localized Re-Reasoning
Traditional frameworks often re-execute the entire workflow when reasoning errors are detected, leading to substantial correction costs and unnecessary recomputation. TALM instead supports localized re-reasoning: only the agents within the affected subtree are recomputed, while the rest of the tree remains intact. This strategy reduces redundant computation, preserves verified reasoning chains, and enhances overall efficiency. By leveraging the hierarchical structure, TALM introduces selective error recovery mechanisms that improve robustness without sacrificing scalability.
Localized re-reasoning in TALM is realized through two complementary modes: clarification-based and structure-correction. These two modes address errors that arise at different stages of the task execution process and together provide a systematic error-handling capability.
@@ 3.3. Long-Term Memory
Most existing multi-agent frameworks are stateless, meaning that agents operate on a one-off basis without retaining knowledge from past experiences. As a result, similar tasks must be repeatedly solved from scratch, and opportunities to leverage prior solutions are lost. To address this limitation, TALM integrates a long-term memory module that enables agents to store, retrieve, and reuse knowledge across tasks.
@@
Figure 3. Long-term memory mechanisms in TALM. The top part illustrates knowledge retrieval, where a new task is encoded and matched with similar past experiences in the vector database. The bottom part illustrates memory update, where validated outputs are stored and merged with existing records when overlap is detected.:
After the code agent completes its subtask and the output passes validation, the memory update mechanism is triggered. The verified code, together with the associated problem description, reasoning process, and metadata, is encoded and stored back into the vector database. To avoid redundancy, newly encoded entries are compared with existing records, and if a highly similar entry is found, a merging strategy is applied to consolidate their contents rather than adding a duplicate record. This iterative update process ensures that the memory remains concise, consistent, and increasingly useful for future reasoning. This update process is illustrated in the bottom part of Figure 3.
By integrating tree-structured collaboration, localized re-reasoning, and long-term memory, TALM balances flexibility, efficiency, and continual improvement, positioning it as a robust framework for complex code generation.
@@ 4.1.2. Baselines
Together, these baselines cover different paradigms—from direct prompting to advanced reasoning and multi-agent collaboration—providing a comprehensive comparison to highlight TALM’s contributions.
@@ 4.1.3. Foundation Models, Evaluation Metric, & Hyperparameters
For the main results reported in Section 4.2, we conduct experiments with both GPT-4o-mini and GPT-4o to ensure a comprehensive comparison across methods. For the ablation studies in Section 4.3, we use GPT-4o-mini only, due to cost considerations. In all experiments, the temperature is fixed at 0 to maintain determinism.
@@ 4.2.2. Token Usage Efficiency
We further compare TALM with and without the long-term memory module. Results show only a marginal increase in token usage when memory is enabled, since retrieval and storage are handled externally by the vector database rather than the LLM itself. This indicates that TALM achieves accuracy gains with almost no additional token cost, making it well suited for large-scale, multi-round program synthesis tasks.

### v2

@@ Abstract.
conference: International Workshop on Agentic Engineering; Apr. 14, 2026; Rio de Janeiro, Brazilccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
We evaluate TALM on HumanEval (Chen et al., 2021), BigCodeBench (Zhuo et al., 2025), and ClassEval (Du et al., 2023) benchmarks. Results show that TALM outperforms existing approaches while maintaining high token efficiency. Beyond benchmark results, TALM provides a general paradigm for integrating structured reasoning with long-term memory, suggesting a path forward for scalable, reliable and self-improving code generation agents.
@@ 2. Related Work
AutoGen (Wu et al., 2023) provides a general multi-agent platform where a centralized planner coordinates multiple specialized agents through role-playing and structured dialogues. This design improves modularity and allows flexible integration of external tools. TDAG (Wang et al., 2025) dynamically decomposes tasks into executable steps and instantiates specialized sub-agents as needed. Its reflection mechanism enables feedback and backtracking, enhancing robustness and recovery. AgentNet (Yang et al., 2025) employs a dynamically evolving DAG communication topology, allowing agents to adapt roles and strategies through an evolution phase that fosters specialization and autonomous task allocation. In contrast, ChatDev (Qian et al., 2024) focuses on simulating the software development life cycle by assigning agents to roles such as analyst, designer, coder, and tester, with dedicated mechanisms for cross-stage context menagement. MetaGPT (Hong et al., 2024) emphasizes responsibility separation: each agent operates within well-defined input–output boundaries, and communication is achieved through a subscription-based messaging mechanism.
@@ 3. TALM Framework
Figure 1. TALM workflow overview. The root agent decomposes the task into a tree of subtasks; child agents solve localized steps and return results upward.
A page-wide diagram showing a root agent splitting a task into a hierarchical tree; local verification triggers subtree-only reruns; logs flow into a memory store.
@@ 3.1. Tree-Structured Collaboration
(a) Child-Agent Clarification (bottom-up).
A child agent requests clarifications from its parent when the task specification is ambiguous; the parent refines requirements, enabling correct implementation.
(b) Structure Correction (top-down).
The parent agent identifies a flawed decomposition and issues a corrected structure; only the affected subtree is recomputed.
Figure 2. Localized re-reasoning in TALM. Left: child-to-parent clarification when task specs are underspecified; Right: parent-initiated structure correction when decomposition is flawed.
Two diagrams: (a) shows bottom-up clarifications; (b) shows top-down correction and targeted subtree recomputation.
To prevent uncontrolled tree growth, TALM introduces three hyperparameters: the maximum depth $m$, the initial branching factor $n$, and the decay rate $k$ ($m,n,k\in\mathbb{Z}^{+}$). A node may create children only if its depth is less than $m$, ensuring that nodes at depth $m$ are leaves. The number of children a node may spawn begins at $n$ and decreases by $k$ at each subsequent level, producing a narrowing tree shape. This decay mechanism encourages deeper nodes to concentrate on highly specific subtasks, while also constraining overall complexity. Together, these parameters regulate the height and width of the collaboration tree, striking a balance between flexibility, resource efficiency, and structural controllability. [^1] The decay mechanism is designed under the assumption that task decomposition progressively reduces problem complexity. In the rare cases where a subtask does not become simpler or may even increase in difficulty during decomposition, the framework still allows the subtask to be handled by a single agent without further branching, relying on the agent’s standalone problem-solving capability.
@@ 3.2. Localized Re-Reasoning
Traditional frameworks often re-execute the entire workflow when reasoning errors are detected, leading to substantial correction costs and unnecessary recomputation. TALM instead supports localized re-reasoning: only the agents within the affected subtree are recomputed, while the rest of the tree remains intact. This strategy reduces redundant computation, preserves verified reasoning chains, and enhances overall efficiency. By leveraging the hierarchical structure, TALM introduces localized error recovery mechanisms that improve robustness without sacrificing scalability.
Localized re-reasoning in TALM is realized through two complementary modes: clarification-based and structure-correction (Figure 2). These two modes address errors that arise at different stages of the task execution process and together provide a systematic error-handling capability.
@@ 3.3. Long-Term Memory
Most existing multi-agent frameworks are stateless, meaning that agents operate on a one-off basis without retaining knowledge from past experiences. As a result, similar tasks must be repeatedly solved from scratch. To address this limitation, TALM incorporates a long-term memory module that allows agents to store, retrieve, and reuse knowledge across tasks, enabling cumulative learning and improved efficiency over time.
@@
Figure 3. Long-term memory mechanisms in TALM. The top part illustrates knowledge retrieval, where a new task is encoded and matched with similar past experiences in the vector database. The bottom part illustrates memory update, where validated outputs are stored and merged with existing records when overlap is detected.
A diagram with two sections: the top shows knowledge retrieval, where a task description is encoded and searched against past records in a vector database; the bottom shows memory update, where newly validated results are added or merged with existing entries to avoid redundancy.
After the code agent completes its subtask and the output passes validation, the memory update mechanism is triggered. The verified code, together with the associated problem description, reasoning process, and metadata, is encoded and stored back into the vector database. Unlike standard retrieval-augmented generation (RAG) approaches (Lewis et al., 2020; Gao et al., 2024) that treat the memory as a static repository for query-time retrieval only, TALM actively maintains and refines its long-term memory through consolidation.
Specifically, the newly encoded entry is compared against existing memory records using vector similarity search. Past entries whose similarity exceeds a predefined threshold (0.95 in our implementation) are retrieved and jointly considered with the new entry. These highly similar records are then provided to LLM, which synthesizes them into a single representative memory entry that unifies their shared knowledge. This synthesis produces updated versions of the problem description, reasoning process, and generated code, capturing the common structure and solution patterns across experiences. The consolidated entry is stored back into the memory, while the original overlapping records are removed. This process prevents uncontrolled memory growth, reduces redundancy, and mitigates context contamination, thereby keeping the memory compact, coherent, and efficient for sustained multi-task reasoning. This update process is illustrated in the bottom part of Figure 3.
By integrating tree-structured collaboration, localized re-reasoning, and long-term memory, TALM balances flexibility, efficiency, and continual improvement, positioning it as a robust framework for complex code generation. A representative code generation example produced by TALM is provided in Appendix A.
@@ 4.1.2. Baselines
Together, these baselines cover different paradigms—from direct prompting to advanced reasoning and multi-agent collaboration—providing a comprehensive comparison to highlight TALM’s contributions. [^2] We note that some well-known multi-agent frameworks, such as MetaGPT and AgentNet, are not included as baselines. MetaGPT is excluded because its outputs are difficult to evaluate automatically in our experimental setup, while AgentNet was released after the experiments in this paper were conducted.
@@ 4.1.3. Foundation Models, Evaluation Metric, & Hyperparameters
For the main results reported in Section 4.2, we conduct experiments with both GPT-4o-mini and GPT-4o to ensure a comprehensive comparison across methods. For the ablation studies in Section 4.3, we use GPT-4o-mini only, due to cost considerations. In all experiments, the temperature is fixed at 0 to maintain determinism. Also, for all baseline approaches, we re-implement and evaluate them using the same foundation models (GPT-4o-mini and GPT-4o), rather than directly adopting results from prior work, to ensure a fair comparison.
Due to budget constraints, we limit our experiments to these two models. Under a fixed budget, we argue that comparing methods across different capability tiers of the same model family provides a clearer assessment of generalization, compared to a horizontal comparison across models from different providers (e.g., OpenAI vs. Anthropic) or across open-source and closed-source models, where architectural and training differences may introduce additional confounding factors.
@@ 4.2.2. Token Usage Efficiency
We further compare TALM with and without the long-term memory module. Results show only a marginal increase in token usage when memory is enabled, since retrieval and storage are handled externally by the vector database rather than the LLM itself. This indicates that TALM achieves accuracy gains with few additional token cost, making it well suited for large-scale, multi-round program synthesis tasks.
@@ 5. Conclusions and Future Work
Our results also reveal a trade-off introduced by task decomposition. For relatively simple tasks, decomposition may introduce additional overhead or occasional failures due to increased token usage and compounding reasoning errors. In contrast, for more challenging problems, decomposition substantially improves solvability by breaking down complexity, effectively lowering the performance floor while raising the ceiling—an effect we consider more impactful in realistic software engineering scenarios.
@@ 6. Limitations
Appendix A Code Generation Sample by TALM
Here we provide a representative code sample generated by TALM while handling a task from BigCodeBench. The example demonstrates TALM’s ability to decompose problems and generate structured, executable code, offering qualitative insight beyond the quantitative evaluation reported in the main paper.
$\begin{small}${
import cgi
import http.server
import smtplib
from email.mime.text import MIMEText
import json
def task_func(smtp_server, smtp_port, smtp_username, smtp_password):
REQUIRED_KEYS = ("subject", "message", "to")
class EmailRequestHandler(http.server.BaseHTTPRequestHandler):
server_version = "EmailHTTP/1.0"
sys_version = ""
def do_POST(self):
try:
content_type = self._require_json_content_type()
body_bytes = self._read_request_body()
email_data = self._parse_json(body_bytes)
self._validate_email_data(email_data)
msg = self._build_mime_message(email_data)
self._send_email(msg, email_data["to"])
self._send_ok()
except json.JSONDecodeError:
self._send_error_response(400, "Bad Request: invalid JSON")
except ValueError as e:
self._send_error_response(400, f"Bad Request: {e}")
except smtplib.SMTPAuthenticationError:
self._send_error_response(535, "Authentication Failed")
except Exception:
self._send_error_response(500, "Internal Server Error")
def _require_json_content_type(self) -> str:
raw = self.headers.get("content-type")
if not raw:
raise ValueError("Missing Content-Type header")
ctype, _pdict = cgi.parse_header(raw)
if ctype.lower() != "application/json":
raise ValueError("Content-Type must be application/json")
return ctype
def _read_request_body(self) -> bytes:
raw_len = self.headers.get("content-length")
if raw_len is None:
raise ValueError("Missing Content-Length header")
try:
length = int(raw_len)
except ValueError:
raise ValueError("Invalid Content-Length header")
if length < 0:
raise ValueError("Invalid Content-Length header")
return self.rfile.read(length)
def _parse_json(self, body: bytes) -> dict:
text = body.decode("utf-8")
data = json.loads(text)
if not isinstance(data, dict):
raise ValueError("JSON body must be an object")
return data
def _validate_email_data(self, data: dict) -> None:
missing = [k for k in REQUIRED_KEYS if k not in data]
if missing:
raise ValueError(f"Missing required key(s): {’, ’.join(missing)}")
if not isinstance(data["subject"], str) or not data["subject"].strip():
raise ValueError("’subject’ must be a non-empty string")
if not isinstance(data["message"], str):
raise ValueError("’message’ must be a string")
if not isinstance(data["to"], str) or "@" not in data["to"]:
# Not full RFC validation; avoids obvious garbage
raise ValueError("’to’ must be a valid email-like string")
def _build_mime_message(self, data: dict) -> MIMEText:
msg = MIMEText(data["message"], _subtype="plain", _charset="utf-8")
msg["Subject"] = data["subject"]
msg["From"] = smtp_username
msg["To"] = data["to"]
return msg
def _send_email(self, msg: MIMEText, recipient: str) -> None:
with smtplib.SMTP(smtp_server, smtp_port, timeout=20) as server:
server.ehlo()
try:
server.starttls()
server.ehlo()
except smtplib.SMTPException:
pass
server.login(smtp_username, smtp_password) # may raise SMTPAuthenticationError
server.sendmail(smtp_username, [recipient], msg.as_string())
def _send_ok(self) -> None:
body = b"OK"
self.send_response(200)
self.send_header("Content-Type", "text/plain; charset=utf-8")
self.send_header("Content-Length", str(len(body)))
self.end_headers()
self.wfile.write(body)
def _send_error_response(self, code: int, message: str) -> None:
body = message.encode("utf-8")
self.send_response(code)
self.send_header("Content-Type", "text/plain; charset=utf-8")
self.send_header("Content-Length", str(len(body)))
self.end_headers()
self.wfile.write(body)
def log_message(self, fmt, *args):
return
return EmailRequestHandler
[…]